OpenStudy (sphott51):

Solve 2(x+3)^2 -4 =6 the answer is x= -3 +- square root of 5 Just need to show the work and idk how.

jagr2713 (jagr2713):

correct :D

jagr2713 (jagr2713):

ok now we are going to work on this :D

jagr2713 (jagr2713):

since -4 does not contain the variable to solve for, move it to the right side of the equation by adding 4 to both sides 2(x+3)^2=4+6

jagr2713 (jagr2713):

add 4 and 6 to get 10 2(x+3)^2=10

jagr2713 (jagr2713):

what do you think we do next @SPHott51

OpenStudy (radar):

Divide thru by 2 getting (x+3)^2 -2 = 3 now add 2 to both sides; That gives you (x+3)^2 = 5 Now take the sqrt of both sides. Giving: x + 3 = sqrt 5 Now solve for x getting you what you expected.

jagr2713 (jagr2713):

@SPHott51 you there

OpenStudy (sphott51):

Yeah, I'm here

jagr2713 (jagr2713):

do you know what we do next :D

OpenStudy (sphott51):

a few.. Just dont know to to solve x-3= sqrt 5 to get the x..

jagr2713 (jagr2713):

well next we divide each term in the equation by 2 2(x+3)^2/2=10/2

OpenStudy (sphott51):

I got it now

jagr2713 (jagr2713):

ok :D

OpenStudy (sphott51):

Thanks!

jagr2713 (jagr2713):

yw